### v1

# AI Foundation Models in Remote Sensing: A Survey

Abstract
Artificial Intelligence (AI) technologies have profoundly transformed the field of remote sensing, revolutionizing data collection, processing, and analysis. Traditionally reliant on manual interpretation and task-specific models, remote sensing has been significantly enhanced by the advent of foundation models—large-scale, pre-trained AI models capable of performing a wide array of tasks with unprecedented accuracy and efficiency. This paper provides a comprehensive survey of foundation models in the remote sensing domain, covering models released between June 2021 and June 2024. We categorize these models based on their applications in computer vision and domain-specific tasks, offering insights into their architectures, pre-training datasets, and methodologies. Through detailed performance comparisons, we highlight emerging trends and the significant advancements achieved by those foundation models. Additionally, we discuss the technical challenges, practical implications, and future research directions, addressing the need for high-quality data, computational resources, and improved model generalization. Our research also finds that pre-training methods, particularly self-supervised learning techniques like contrastive learning and masked autoencoders, significantly enhance the performance and robustness of foundation models in remote sensing tasks such as scene classification, object detection, and other applications. This survey aims to serve as a resource for researchers and practitioners by providing a panorama of advances and promising pathways for continued development and application of foundation models in remote sensing.
Index Terms:
Remote sensing, Machine learning, Artificial intelligence, Image processing, Computer vision, Transformers.
I Introduction
Artificial Intelligence (AI) technologies have profoundly transformed the field of remote sensing, revolutionizing how data is collected, processed, and analyzed. Traditionally, remote sensing projects relied heavily on manual interpretation and task-specific models that required extensive labeled datasets and significant computational resources. However, the advent of AI and deep learning (DL) has ushered in a new era in which large-scale, pre-trained models, known as foundation models, are capable of performing a wide array of tasks with unprecedented accuracy and efficiency. These advancements have not only enhanced the capabilities of remote sensing but have also opened new avenues for its application across various domains.
In recent years, numerous foundation models have emerged, demonstrating remarkable performance in handling diverse remote sensing tasks. These models have shown the potential to significantly improve performance on multiple downstream tasks such as scene classification, semantic segmentation, object detection, and more. By leveraging vast amounts of pre-training data and sophisticated architectures, these foundation models have set new benchmarks in the field, making them indispensable tools for researchers and engineers alike.
This paper aims to provide a comprehensive survey of foundation models in the remote sensing domain, covering foundation models released between June 2021 and June 2024. In figure 1, 51 vision models are listed in chronological order. To facilitate easy navigation and utility for researchers, we have categorized these models based on their application in computer vision tasks and domain-specific tasks. This categorization allows for a clearer understanding of which models are suited for certain purposes, whether they involve general image-based challenges or more specialized applications such as environmental monitoring, land cover and land use, archaeological prospection, disaster management, or other domains.
Our contributions include:
  1. 1.
    An exhaustive review of current state of foundation models proposed in the field of remote sensing, starting from the background and methodologies of foundation models to specific applications across different domains and tasks in a hierarchical and structured manner.
  2. 2.
    Categorization and analysis of the models based on their application in both computer vision tasks (table I) and domain-specific tasks (table 5). We discuss the architecture, pre-training datasets, pre-training methods, and performances of each model.
  3. 3.
    Discussion of challenges and unresolved aspects related to foundation models in remote sensing. We pinpoint new trends, raise important questions, and proposed future directions for further exploration.
Figure 1: Overview of some well-known foundation models for remote sensing from 2021 June to 2024 June.
Figure 2: Examples of data types used in those foundation models and downstream tasks that can be done by foundation models. Data: (1) Panchromatic[2], (3) True Color, (4) SAR[68], (5) Hyperspacial[2], (6) Multispectral[2]. Downstream tasks: (1) Segmentation, (2) Object Detection, (3) Classification[10], (4) Change Detection[53].
 [^1] True Color, Segmentation, and Object detection images © MAXAR 2024, provided through the NextView License Agreement.
II Background
II-A Remote Sensing
Remote sensing (RS) refers to the process of acquiring information about objects or areas from a distance, typically using satellite or airborne sensors. These technologies and techniques serve vital roles in diverse fields, enabling the collection of data over land geographic areas without physical contact. Applications of remote sensing include earth observation, digital archaeology, urban planning and development, and disaster management. The field of remote sensing has been developed rapidly since the mid 20th century. Initially, remote sensing relied on analog photographic techniques via aerial and satellite platforms, which provided limited spectral and spatial resolution. The launch of early Earth observation satellites, such as Landsat program commenced in 1967 [78], marked a significant advancement, enabling consistent and wide-ranging data collection for environmental monitoring.
Modern remote sensing employs a variety of sensors, including optical, thermal, and radar, each suited for specific types of data collection. Optical sensors capture a wide variety of spectral bands, including visible and near-infrared light, allowing for detailed imaging of land cover and vegetation health. Thermal sensors detect heat emitted or reflected from the Earth’s surface, useful for monitoring volcanic activity, forest fires, and climate change monitoring. Radar sensors can penetrate clouds and vegetation, providing critical information in all-weather conditions and for applications such as soil moisture estimation and urban infrastructure mapping [50, 12].
Year-Month Architecture Model Name Scene Classification Semantic Segmentation Object Detection Change Detection
2021 Jun ResNet-50 CMC-RSSR [59] $\checkmark$
2021 Oct ResNet-50 SeCo [46] $\checkmark$ $\checkmark$
2021 Oct ResNet-50 GeoKR [39] $\checkmark$ $\checkmark$ $\checkmark$
2021 Dec ResNet-34 MATTER [1] $\checkmark$ $\checkmark$ $\checkmark$
2022 Mar ResNet-50 GASSL [3] $\checkmark$ $\checkmark$ $\checkmark$
2022 May ViTAEv2-S RSP [69] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2022 Jun ViT-S/8 DINO-MM [73] $\checkmark$
2022 Jun Swin Transformer Scheibenreif, et al. [56] $\checkmark$ $\checkmark$
2022 Jul ViT/Swin Transformer RingMo [61] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2022 Aug ResNet-50 GeCO [40] $\checkmark$ $\checkmark$ $\checkmark$
2022 Sep BYOL RS-BYOL [32] $\checkmark$ $\checkmark$
2022 Nov ViT-B CSPT [87] $\checkmark$ $\checkmark$
2022 Nov ViT RVSA [71] $\checkmark$ $\checkmark$ $\checkmark$
2023 Jan MAE-based Framework SatMAE [11] $\checkmark$ $\checkmark$
2023 Apr TOV TOV [63] $\checkmark$ $\checkmark$ $\checkmark$
2023 Apr
Teacher-student
Self-distillation
CMID [49] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Jun CACo CACo [47] $\checkmark$ $\checkmark$ $\checkmark$
2023 Jun ResNet-18 IaI-SimCLR [54] $\checkmark$
2023 Aug Teacher-Student GFM [48] $\checkmark$ $\checkmark$ $\checkmark$
2023 Aug Swim Transformer SatLasPretrain [4] $\checkmark$ $\checkmark$
2023 Sep Multi-Branch RingMo-Sense [85] $\checkmark$
2023 Sep ViT Scale-MAE [55] $\checkmark$ $\checkmark$
2023 Sep CNN-Transformer RingMo-lite [76] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Sep Multimodel SSL DeCUR [72] $\checkmark$ $\checkmark$
2023 Oct MSFE+MMFH Feng et al. [21] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Oct ViT FG-MAE [75] $\checkmark$ $\checkmark$
2023 Nov ViT Prithvi [33] $\checkmark$
2023 Nov Multimodal Encoder CROMA [22] $\checkmark$ $\checkmark$
2023 Dec ViT USat [31] $\checkmark$
2024 Jan ViT-B Cross-Scale MAE [62] $\checkmark$ $\checkmark$
2024 Jan Unet+Transformer U-BARN [20] $\checkmark$ $\checkmark$
2024 Jan
Autoregressive
Transformer
EarthPT [58] $\checkmark$
2024 Jan Teacher-Student Network GeRSP [29] $\checkmark$ $\checkmark$ $\checkmark$
2024 Jan Dual-Branch SwiMDiff [65] $\checkmark$ $\checkmark$
2024 Jan Generative ConvNet SMLFR [17] $\checkmark$ $\checkmark$
2024 Feb 3D GPT SpectralGPT [28] $\checkmark$ $\checkmark$ $\checkmark$
2024 Feb MAE-based Framework Presto [66] $\checkmark$
2024 Mar SatMAE SatMAE++ [52] $\checkmark$
2024 Mar
Joint-Embedding
Predictive Architecture
SAR-JEPA [41] $\checkmark$
2024 Mar ViT FoMo-Bench [5] $\checkmark$ $\checkmark$ $\checkmark$
2024 Mar
Factorized Multi-Modal
Spatiotemporal Encoder
SkySense [24] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 Mar Multi-Modules UPetu [18] $\checkmark$ $\checkmark$ $\checkmark$
2024 Apr Swim Transformer msGFM [25] $\checkmark$ $\checkmark$
2024 Apr DINO DINO-MC [77] $\checkmark$ $\checkmark$
2024 May OFA-Net OFA-Net [84] $\checkmark$ $\checkmark$
2024 May
Shared Encoder,
Task-Specific Decoders
MTP [70] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 May ViT BFM [7] $\checkmark$ $\checkmark$
2024 May MP-MAE MMEarth [51] $\checkmark$ $\checkmark$
2024 May ViT CtxMIM [86] $\checkmark$ $\checkmark$ $\checkmark$
2024 May HiViT SARATR-X [37] $\checkmark$ $\checkmark$
2024 Jun Dynamic OFA DOFA [83] $\checkmark$
TABLE I: Summery of computer-vision downstream tasks tested by each models.
In recent years, remote sensing has found applications in many fields. In environmental monitoring, it is used to track deforestation, to monitor air and water quality, and to assess the impacts of climate change [23]. In agriculture, remote sensing helps in crop health monitoring, yield estimation, and efficient resource management [50]. Urban planning and development benefit from remote sensing through the monitoring of urban sprawl, infrastructure development, and land-use planning [12]. Furthermore, in disaster management, remote sensing is crucial for assessing the damage caused by natural disasters, aiding in the planning and execution of relief operations [23].
The integration of remote sensing data with Geographic Information Systems (GIS) has further enhanced its utility. GIS provides a framework for capturing, storing, analyzing, and visualizing spatial and geographic data. When combined with remote sensing data, GIS can be used to create detailed and dynamic maps and models for various applications. This synergy is particularly valuable in resource management, urban planning, and disaster response, where accurate and timely information is critical [50, 12, 23].
II-B Foundation Models for Remote Sensing
Foundation models (FMs) refer to large-scale, pre-trained models that provide a robust starting point for various downstream tasks across different domains [34]. These models leverage extensive datasets and advanced architectures, enabling them to capture complex patterns and features that can be fine-tuned for specific applications with minimal additional training. In remote sensing, FMs are particularly valuable due to the diverse and complex nature of the data, including multi-spectral and multi-temporal imagery. Techniques such as self-supervised learning (SSL) [35] and transformers [67] have significantly enhanced the performance and efficiency of tasks such as image classification, object detection, and change detection, addressing the unique challenges posed by remote sensing data [14].
The evolution of FMs has been driven by advancements in deep learning and the availability of large datasets. Initially, convolutional neural networks (CNNs) like ResNet [26] paved the way for improved image recognition and classification tasks [45]. The introduction of transformers, which use self-attention mechanisms to model long-range dependencies, has further advanced the capabilities of FMs in handling large-scale image data [11].
FMs in remote sensing are characterized by their ability to leverage vast amounts of unlabeled data through SSL techniques, allowing them to learn robust representations without the need for extensive labeled datasets [88]. Key SSL methods include contrastive learning, which involves learning representations by comparing different augmented views of the same data point, and predictive coding, which trains models to predict missing parts of the input data from the observed parts [55].
Notable foundation models in remote sensing include SatMAE [11], which pre-trains transformers for temporal and multi-spectral satellite imagery; Scale-MAE[55], a scale-aware masked autoencoder for multiscale geospatial representation learning ; and DINO-MC [77], which extends global-local view alignment for SSL with remote sensing imagery. These models have shown remarkable performance in various remote sensing tasks such as scene classification, object detection, and change detection.
Despite their success, FMs face several challenges, including the need for high-quality and diverse training data, significant computational resources, and effective domain adaptation to specific remote sensing tasks [52]. Addressing these challenges will be crucial for the continued advancement of FMs in remote sensing.
III Methodologies
In recent years, the development of foundation models (FMs) for remote sensing has leveraged various sophisticated methodologies. These include self-supervised learning (SSL), transformers and vision transformers (ViT), and residual neural networks (ResNet). These methodologies have significantly enhanced the capabilities of FMs by enabling them to learn from large amounts of data without extensive human supervision, handle complex data structures, and improve feature extraction and representation. This section reviews the mechanisms and contributions of these methodologies to the field of remote sensing.
III-A Self-Supervised Learning (SSL) Training Strategies for Remote Sensing FMs
Self-supervised learning (SSL) plays a crucial role in the pre-training stage of a foundation model. Through SSL, a model learns to predict part of the input from other parts of the input data, reducing the reliance on large labeled datasets. SSL is particularly valuable in remote sensing due to the scarcity of labeled data. Models pre-trained with SSL can effectively capture patterns and features from vast amounts of unlabeled remote sensing data, making them highly efficient for downstream tasks. Figure 2 presents a general pipeline of self-supervised learning.
Figure 3: General pipeline of SSL[35]. Diverse datasets images and pretext task images are acquired from ImageNet[13], BigEarthNet[60], and MillionAID[44]. Finetune dataset includes images from DIOR[38].
 [^2] Object detection and Segmentation © MAXAR 2024, provided through the NextView License Agreement.
In remote sensing applications, two specific SSL methods used commonly include contrastive learning and predictive coding.
III-A1 Contrastive Learning.
Contrastive learning aims to learn representations by comparing different augmented views of the same data point. The main idea is to bring similar (positive) pairs closer and push dissimilar (negative) pairs apart in the feature space. This method relies heavily on data augmentation to create multiple views of the same image.
III-A2 Predictive Coding.
Predictive coding is another SSL technique where the model learns to predict missing parts of the input data from the observed parts. This method is beneficial for capturing spatial and temporal dependencies within the data. Common predictive coding methods include autoencoders (AE) and masked autoencoders (MAE).
Common SSL methods include SimCLR, MoCo (Momentum Contrast), BYOL (Bootstrap Your Own Latent), and DINO (Self-Distillation with No Labels). Each of these methods has a distinctive approach to creating positive and negative pairs and updating model parameters. These methodologies have proven effective in tasks such as scene classification, semantic segmentation, and object detection in remote sensing. For example, SSL can help models learn to classify land cover types, identify objects like buildings and vehicles, and segment different regions within satellite images, even when labeled data is limited.
III-B Backbones
In deep learning, backbones are essential neural network architectures that serve as feature extractors. They form the foundational layers of a model, processing input data to generate rich, hierarchical feature representations. These representations can then be utilized by subsequent model components to perform various tasks such as classification, detection, and segmentation. By leveraging powerful backbones, models can efficiently handle complex data and improve their performance across different applications. In the following section, we
III-B1 Backbone Type I: Transformers and Vision Transformers (ViT)
Transformers, originally designed for natural language processing, have revolutionized the field of computer vision with their ability to model long-range dependencies through self-attention mechanisms. Vision Transformers (ViT) [19] adapt the transformer architecture proposed by Vaswani et al. (2017)[67] to handle image data by treating image patches as sequences of tokens. This adaptation is particularly useful in remote sensing, where images can be large and contain complex spatial structures. Figure 3 illustrates the basic structure of ViT for RS segmentation tasks.
The key components of ViT include patch embedding, positional encoding, the transformer encoder, and the classification head [19]. Patch embedding involves dividing an image into fixed-size patches and linearly embedding each patch into a vector. Positional encoding adds positional information to the patch embeddings to retain the spatial structure of the image. The transformer encoder consists of multiple layers of multi-head self-attention and feed-forward neural networks, which process the sequence of embedded patches to capture global dependencies. Finally, the classification head is a fully connected layer that processes the final sequence representation for downstream tasks such as image classification. The self-attention mechanism in transformers allows each token to attend to all other tokens, providing a powerful way to capture global context. The attention weights are computed as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
where $Q$ (query), $K$ (key), and $V$ (value) are the input matrices, and $d_{k}$ is the dimension of the key vectors [67]. ViTs have shown great promise in remote sensing tasks such as land cover classification, urban area identification, and vegetation analysis, leveraging their ability to capture both local and global patterns.
Figure 4: The Vision transformer architecture.
 [^3] RGB Image and Segmentation © MAXAR 2024, provided through the NextView License Agreement.
III-B2 Backbone Type II: Convolutional Neural Network (CNN)
Convolutional Neural Networks (CNN), such as Residual Neural Networks (ResNet), address the degradation problem in deep neural networks by introducing residual connections, which allow gradients to bypass certain layers, facilitating the training of very deep networks. This capability is particularly beneficial in remote sensing, where deep models are often required to capture the intricate details and variations in satellite images.
ResNet, as an example, is characterized by their residual blocks, which include shortcut connections that bypass one or more layers. The residual block can be described by the following equation:
$\mathbf{y}=\mathcal{F}(\mathbf{x},\{W_{i}\})+\mathbf{x}$
where $\mathbf{y}$ is the output, $\mathcal{F}$ represents the residual mapping to be learned, $\mathbf{x}$ is the input, and $\{W_{i}\}$ are the weights of the layers in the block[26]. Depending on the dimensions, shortcuts can be identity mappings (if input and output dimensions match) or convolutional layers (if dimensions differ).
ResNet has various architectures like ResNet-50, ResNet-101, and ResNet-152, with the number indicating the total layers. These networks have shown remarkable performance in various vision tasks due to their ability to train deeper networks without degradation. In remote sensing, ResNets are widely used for image classification, object detection, and change detection tasks. For example, ResNet-based models can classify different land cover types, detect objects like buildings and vehicles, and monitor changes in the landscape over time by comparing temporal sequences of satellite images.
By incorporating these methodologies, foundation models for remote sensing can leverage vast amounts of data, handle complex structures, and achieve state-of-the-art performance across various applications. These methodologies enable models to effectively address the unique challenges of remote sensing, such as large image sizes, diverse data sources, and the need for high accuracy in environmental monitoring and analysis.
In the following sections, we will explore specific applications of these methodologies in different remote sensing tasks, analyze their performance, and discuss the datasets used to train and evaluate these models.
IV Datasets
Datasets play a crucial role in remote sensing, providing the foundation for training and evaluating models. High-quality datasets enable models to learn accurate representations of the Earth’s surface, improving their performance on various remote sensing tasks. In this section, we provide an overview of commonly used datasets in table II for remote sensing, discussing their characteristics, applications, and relevance to foundation models.
The RSD46-WHU [42, 82] dataset, introduced in 2017, is sourced from Google Earth and Tianditu. It contains 117,000 images with a patch size of 256 pixels and spatial resolutions ranging from 0.5 to 2 meters per pixel. Covering 46 categories globally, this dataset is primarily used for scene classification. Similarly, the Functional Map of the World (fMoW) [10], released in April 2018, comprises over 1 million images from Digital Globe. Spanning 63 categories across 207 countries, it includes multispectral images used for both scene classification and object detection.
In May 2019, the DOTA [80] dataset was proposed, known for its large-scale aerial image object detection capabilities. It includes 11,268 images of various resolutions from Google Earth, GF-2 Satellite, and aerial sources, covering 18 categories globally. Another significant dataset, SEN12MS [57], released in June 2019, contains 541,986 images from Sentinel-1, Sentinel-2, and MODIS Land Cover. With a patch size of 256x256 pixels, it supports land cover classification and change detection tasks.
BigEarthNet [60], also from June 2019, consists of 590,326 images with varying sizes from 20x20 to 120x120 pixels, sourced from Sentinel-2. It covers 43 categories across Europe and is used for scene classification and object detection. The SeCo [46] dataset, another June 2019 release, contains approximately 1 million images with a resolution of 2.65x2.65km from Sentinel-2. It is designed for seasonal change detection and land cover classification over seasons.
The MillionAID dataset [43, 44], introduced in March 2021, includes over 1 million images of various sizes from Google Earth. Covering 51 categories globally, it is used for scene classification. Levir-KR, released in July 2021, contains 1,431,950 images from Gaofen-1, Gaofen-2, and Gaofen-6 satellites, supporting change detection and scene classification applications.
SoundingEarth [27], introduced in August 2021, comprises 50,545 images of 1024-pixel size from Google Earth, combining RGB and audio data for remote sensing. The TOV-RS-Balanced dataset [64] from April 2022 includes 500,000 images with a 600-pixel size from Google Earth, covering 31 categories globally, and is used for scene classification, object detection, and semantic segmentation.
SeasoNet [36], released in July 2022, features 1,759,830 images from Sentinel-2 with patch size from 20 to 120 pixels, supporting seasonal scene classification, segmentation, and retrieval over Germany. Lastly, the SSL4EO-S12 dataset [74] from November 2022 contains over 3 million images from Sentinel-1 and Sentinel-2, with patch size of 264x264 pixels. It is used for scene classification and object detection.
These datasets, with their varying resolutions, categories, and geographic coverage, provide a rich resource for advancing remote sensing research and applications. They facilitate the development of robust models capable of addressing diverse challenges in understanding and interpreting Earth’s surface through remote sensing technologies.
! Month, Year Dataset Title Patch Size Size Resolution (m) Source Categories Geographic Coverage Image Type Application 2017 RSD46-WHU [42, 82] - 256 x 256 117,000 0.5 - 2 Google Earth, Tianditu 46 Global RGB Scene Classification Apr, 2018 fMoW [10] Functional Map of the World - 1,047,691 - Digital Globe 63 207 of 247 countries Multispectral Scene Classification, Object Detection May, 2019 DOTA [80] DOTA: A Large-scale Dataset for Object Detection in Aerial Images 800 × 800 to 20,000 × 20,000 11,268 Various Google Earth, GF-2 Satellite, and aerial images 18 Global RGB Object Detection Jun, 2019 SEN12MS [57] SEN12MS – A Curated Dataset of Georeferenced Multi-Spectral Sentinel-1/2 Imagery for Deep Learning and Data Fusion 256 x 256 541,986 10 Sentinel-1, Sentinel-2, MODIS Land Cover 33 Globally distributed SAR/Multispectral Land Cover Classification, Change Detection, Jun, 2019 BigEarthNet [60] BigEarthNet: A Large-Scale Benchmark Archive For Remote Sensing Image Understanding 20 x 20 to 120 x 120 590,326 Various Sentinel-2 43 Europe Multispectral Scene Classification, Object Detection Jun, 2019 SeCo [46] Seasonal Contrast: Unsupervised Pre-Training from Uncurated Remote Sensing Data 264 x 264 $\sim$1M 10 - 60 Sentinel-2 - Global Multispectral Seasonal Change Detection, Land Cover Classification over Seasons Mar, 2021 MillionAID [44, 43] Million-AID 110 - 31,672 1,000,848 Various Google Earch 51 Global RGB Scene Classification Jul, 2021 Levir-KR [39] Geographical Knowledge-driven Representation Learning for Remote Sensing Images - 1,431,950 Various Gaofen1, Gaofen-2, Gaofen-6 8 Global Multispectral Change Detection, Scene Classification Apr, 2022 TOV-RS-Balanced [64] TOV: The Original Vision Model for Optical Remote Sensing Image Understanding via Self-supervised Learning 600 x 600 500,000 1 - 20 Google Earth 31 Global RGB Scene Classification, Object Detection, Semantic Segmentation Jul, 2022 SeasoNet [36] SeasoNet: A Seasonal Scene Classification, Segmentation and Retrieval dataset for satellite Imagery over Germany up to 120 x 120 1,759,830 10 - 60 Sentinel-2 33 Germany Multispectral Scene Classification, Scene Segmentation Nov, 2022 SSL4EO-S12 [74] SSL4EO-S12: A Large-Scale Multi-Modal, Multi-Temporal Dataset for Self-Supervised Learning in Earth Observation 264 x 264 3,012,948 10 - 60 Sentinel-1, Sentinel-2 - Global SAR/Multispectral Scene Classification, Object Detection
TABLE II: This table summarizes a set of most commonly used pre-trained datasets for remote sensing.
V Foundation Models for Remote Sensing
In this section, we provide a comprehensive survey of vision foundation models used in remote sensing. These models are categorized into two groups based on their applications: computer-vision tasks and domain-specific tasks. Each category is further divided into specific tasks to highlight the applications and effectiveness of these models in various remote sensing scenarios. Vision models in remote sensing are designed to process and analyze image data to perform various tasks such as scene classification, semantic segmentation, object detection, and more. Below, we discuss these tasks in detail and provide an overview of relevant models.
V-A Computer Vision Tasks
Computer vision tasks involve the automatic analysis and interpretation of imagery to extract meaningful information. In the context of remote sensing, the primary tasks include scene classification, semantic segmentation, object detection, change detection, and land use/land cover classification. Each of these tasks requires different model architectures and training methodologies to achieve optimal performance. The following sections provide an overview of various models used for these tasks, along with their specific applications and performance metrics. Table I summarizes those downstream tasks tested by each vision models.
  1. 1.
    Scene Classification:
    Scene classification in remote sensing involves categorizing images or parts of images into predefined classes, such as urban, forest, water bodies, and agricultural areas. This task is crucial for land use and land cover mapping, environmental monitoring, and urban planning. Accurate scene classification enables better resource management and policy-making by providing detailed and up-to-date information about land cover and land use patterns. In table III, we provide an overview of various foundation models used for scene classification in remote sensing.
  2. 2.
    Semantic Segmentation:
    Semantic segmentation is a crucial task in the field remote sensing that involves classifying each pixel in an image into a predefined category. This task is crucial for creating detailed maps that show the distribution of different land cover types, such as forests, water bodies, urban areas, and agricultural fields. The high-resolution, pixel-level information provided by semantic segmentation is essential for applications in environmental monitoring, urban planning, disaster management, and precision agriculture. In table IV, we provide an overview of various foundation models used for semantic segmentation in remote sensing.
  3. 3.
    Object Detection:
    Object detection is a critical task in remote sensing that involves identifying and locating objects within an image. Unlike scene classification and semantic segmentation, which focus on classifying entire images or individual pixels, object detection aims to recognize specific objects, such as vehicles, buildings, ships, and other structures, and determine their positions within the image. This task is essential for applications in urban planning, environmental monitoring, disaster response, and security. Table V summarizes all foundation models that can be used for remote sensing object detection tasks.
  4. 4.
    Change Detection:
    Change detection in remote sensing involves identifying and analyzing changes in the Earth’s surface over time by comparing images taken at different times. This task is vital for monitoring environmental changes, urban expansion, deforestation, natural disasters, and other dynamic phenomena. By detecting changes accurately, change detection provides crucial insights for environmental management, urban planning, disaster response, and climate change studies. Foundation models that can accomplish change detection tasks are listed in table VI.
    ! Year-Month Architecture Model Name Pre-training Dataset Resolution (m) Geographic Coverage Visual Encoder Pretrain Methods # of Params 2021 Oct ResNet-50 GeoKR [39] Levir-KR [39] 0.8 to 16 Global ResNet-50/VGG16 Geographical Knowledge Supervision 23.5M/138M 2021 Oct ResNet-50 SeCo [46] Sentinel-2 Imagery 10, 20, and 60 200k Locations Worldwide ResNet-50 CL 23.5M 2022 Jun ViT-S/8 DINO-MM [73] BigEarthNet - MM 10 Global ViT-S/8 SSL 22M 2023 Jan MAE-based Framework SatMAE [11] fMoW Sentinal-2 10, 20, 60 Global ViT-Large MAE 307M 2022 Sep BYOL RS-BYOL [32] Sen12MS [57] 10 to 20 Global ResNet-50 SSL 23.5M 2022 Aug ResNet-50 GeCO [40] Levir-KR [39] 0.8 to 16 Global ResNet-50 SSL 23.5M 2022 Jul ViT/Swin Transformer RingMo [61] 2 million RS images 0.3 to 30 6 Continents ViT/Swin Transformer MIM - 2022 Nov ViT RVSA [71] MillionAID [44, 43] 0.5 to 153 Global ViT-B/ViTAE-B MAE 100M 2021 Dec ResNet-34 MATTER [1] Sentinel-2 Imagery - Rural and Remote Regions with Little Changes ResNet-34 SSL 21.3M 2022 Nov ViT-B CSPT [87] ImageNet-1K - Global ViT-B SSL 86M 2022 Jun Swin Transformer Scheibenreif, et al. [56] SEN12MS 10 Global Swim Transformer CL - 2023 Apr TOV TOV [63] TOV-NI, TOV-RS - Global ViT SSL - 2023 Apr Teacher-student Self-distillation CMID [49] MillionAID [44, 43] Varied Global ResNet-50/Swin-B SSL 25.6M/87.8M 2023 Jun ResNet-18 IaI-SimCLR [54] SEN12MS - Global ResNet-18 CL 11.7M 2023 Jun CACo CACo [47] Sentinel-2 Imagery 10 Global ResNet-18/50 SSL 11.7M/23.5M 2023 Aug Swim Transformer SatLasPretrain [4] SatLasPretrain 1, 10 All Continents except Antarctica Swim Transformer Pretrained on Dataset 88M 2023 Aug Teacher-Student GFM [48] GeoPile - Global Swim Transformer Continual Pretraining - 2023 Sep ViT Scale-MAE [55] FMoW [10] - Global ViT-Large MAE 322.9M 2024 Apr DINO DINO-MC [77] SeCo-100K [46] 10 to 60 Global ViT, Swim Transformer, ResNet, WRN SSL - 2023 Nov Multimodal Encoder CROMA [22] SSL4EO [74] 10 Areas Surrounding Human Settlements ViT-B CL, MAE 86M 2024 Jan ViT-B Cross-Scale MAE [62] fMoW [10] - Global ViT-B MAE 86M 2023 Sep Multimodel SSL DeCUR [72] SSL4EO-S12, RGB-DEM/depth Varied Global ResNet-50 SSL 23.5M 2024 May ViT CtxMIM [86] WorldView-3 Imagery Varied Asia Swim-B MIM 88M 2023 Oct ViT FG-MAE [75] SSL4EO-S12 [74] 10 Global ViT MAE - 2023 Sep CNN-Transformer RingMo-lite [76] AID [81] 0.3 to 30 Global Swim Transformer + CNN FD-MIM 60% less than RingMo 2023 Oct MSFE+MMFH Feng et al. [21] Multi-modal Dataset Varied Global Transformer SSL - 2024 Jan ViT USat [31] Satlas Varied Global ViT MAE - 2024 Mar ViT FoMo-Bench [5] Multiple Varied Global ViT MAE 101M/110M 2024 Jan Unet+Transformer U-BARN [20] Sentinel-2 Imagery Varied France Unet+Transformer SSL - 2024 Jan Teacher-Student Network GeRSP [29] ImageNet[13], MillionAID[44, 43] 0.5 to 153 Global ResNet-50 SSL, SL - 2024 Jan Dual-Branch SwiMDiff [65] Sen12MS [57] Varied Global ResNet-18 SSL 11.7M 2024 May OFA-Net OFA-Net [84] Multi-modal Dataset Varied Global ViT MIM - 2024 Feb 3D GPT SpectralGPT [28] Sentinel-2 Imagery Varied Global ViT MAE 100M/300M/600M 2024 May HiViT SARATR-X [37] SAR Datasets 0.1 to 3 Global HiViT MIM 66M 2024 Mar SatMAE SatMAE++ [52] fMoW [10] Varied Global ViT-Large Multi-Scale Pre-training - 2024 Apr Swim Transformer msGFM [25] GeoPile-2 0.1 to 153 Global Swim Transformer MIM 89M 2024 Mar Factorized Multi-Modal Spatiotemporal Encoder SkySense [24] Multiple Varied Global Multiple CL 2.06B 2024 May Shared Encoder Task-Specific Decoders MTP [70] SAMRS Varied Global RVSA, InternImage Multi-Task Pretraining over 300M 2024 Jun Dynamic OFA DOFA [83] Multiple Varied Global ViT MIM - 2024 Mar Multi-Modules UPetu [18] GeoSense - Global ConvNeXt-B SSL 0.65M 2024 May MP-MAE MMEarth [51] Multi-modal, Geospatial Data - Global ConvNeXt V2 MP-MAE 3.7M to 650M
    TABLE III: Summary of foundation models for scene classification in remote sensing, detailing architecture, pre-training datasets, resolution, geographic coverage, visual encoders, pretraining methods, and number of parameters.
    ! Year-Month Architecture Model Name Pre-training Dataset Resolution (m) Geographic Coverage Visual Encoder Pretrain methods # of Params 2021 Oct ResNet-50 GeoKR [39] Levir-KR [39] 0.8 to 16 Global ResNet-50/VGG16 Geographical Knowledge Supervision 23.5M/138M 2021 Oct ResNet-50 SeCo [46] Sentinel-2 Imagery 10, 20, 60 200k Locations Worldwide ResNet-50 CL 23.5M 2022 Jun ViT-S/8 DINO-MM [73] BigEarthNet - MM 10 Global ViT-S/8 SSL 22M 2023 Jan MAE-based Framework SatMAE [11] fMoW Sentinal-2 10, 20, 60 Global ViT-Large MAE 307M 2022 Sep BYOL RS-BYOL [32] Sen12MS [57] 10 to 20 Global ResNet-50 SSL 23.5M 2022 Aug ResNet-50 GeCO [40] Levir-KR [39] 0.8 to 16 Global ResNet-50 SSL 23.5M 2022 Jul ViT/Swin Transformer RingMo [61] 2 million RS images 0.3 to 30 6 Continents ViT/Swin Transformer MIM - 2022 Nov ViT RVSA [71] MillionAID [44, 43] 0.5 to 153 Global ViT-B/ViTAE-B MAE 100M 2021 Dec ResNet-34 MATTER [1] Sentinel-2 Imagery - Rural and Remote Regions with Little Changes ResNet-34 SSL 21.3M 2022 Nov ViT-B CSPT [87] ImageNet-1K - Global ViT-B SSL 86M 2022 Jun Swin Transformer Scheibenreif, et al. [56] SEN12MS 10 Global Swim Transformer CL - 2023 Apr TOV TOV [63] TOV-NI, TOV-RS - Global ViT SSL - 2023 Apr Teacher-student Self-distillation CMID [49] MillionAID [44, 43] Varied Global ResNet-50/Swin-B SSL 25.6M/87.8M 2023 Jun ResNet-18 IaI-SimCLR [54] SEN12MS - Global ResNet-18 CL 11.7M 2023 Jun CACo CACo [47] Sentinel-2 Imagery 10 Global ResNet-18/50 SSL 11.7M/23.5M 2023 Aug Swim Transformer SatLasPretrain [4] SatLasPretrain 1, 10 All Continents except Antarctica Swim Transformer Pretrained on Dataset - 2023 Aug Teacher-Student GFM [48] GeoPile - Global Swim Transformer Continual Pretraining - 2023 Sep ViT Scale-MAE [55] fMoW [10] - Global ViT-Large MAE 322.9M 2024 Apr DINO DINO-MC [77] SeCo-100K [46] 10 to 60 Global ViT, Swim Transformer, ResNet, WRN SSL - 2023 Nov Multimodal Encoder CROMA [22] SSL4EO [74] 10 Areas Surrounding Human Settlements ViT-B CL, MAE 86M 2024 Jan ViT-B Cross-Scale MAE [62] fMoW [10] - Global ViT-B MAE 86M 2023 Sep Multimodel SSL DeCUR [72] SSL4EO-S12, RGB-DEM/depth Varied Global ResNet-50 SSL 23.5M 2024 May ViT CtxMIM [86] WorldView-3 Imagery Varied Asia Swim-B MIM 88M 2023 Oct ViT FG-MAE [75] SSL4EO-S12 [74] 10 Global ViT MAE - 2023 Oct ViT Prithvi [33] Harmonized Landsat Sentinel 2 30 Contiguous U.S. ViT MAE 100M 2023 Sep CNN-Transformer RingMo-lite [76] AID [81] 0.3 to 30 Global Swim Transformer + CNN FD-MIM 60% less than RingMo 2023 Oct MSFE+MMFH Feng et al. [21] Multi-modal Dataset Varied Global Transformer SSL - 2024 Jan ViT USat [31] Satlas Varied Global ViT MAE - 2024 Mar ViT FoMo-Bench [5] Multiple Varied Global ViT MAE 101M/110M 2024 Jan Unet+Transformer U-BARN [20] Sentinel-2 Imagery Varied France Unet+Transformer SSL - 2024 Jan Teacher-Student Network GeRSP [29] ImageNet[13], MillionAID[44, 43] 0.5 to 153 Global ResNet-50 SSL, SL - 2024 Jan Dual-Branch SwiMDiff [65] Sen12MS [57] Varied Global ResNet-18 SSL 11.7M 2024 May OFA-Net OFA-Net [84] Multi-modal Dataset Varied Global ViT MIM - 2024 Feb 3D GPT SpectralGPT [28] Sentinel-2 Imagery Varied Global ViT MAE 100M/300M/600M 2024 May HiViT SARATR-X [37] SAR Datasets 0.1 to 3 Global HiViT MIM 66M 2024 Mar SatMAE SatMAE++ [52] fMoW [10] Varied Global ViT-Large Multi-Scale Pre-training - 2024 Apr Swim Transformer msGFM [25] GeoPile-2 0.1 to 153 Global Swim Transformer MIM 89M 2024 Mar Factorized Multi-Modal Spatiotemporal Encoder SkySense [24] Multiple Varied Global Multiple CL 2.06B 2024 May Shared Encoder Task-Specific Decoders MTP [70] SAMRS Varied Global RVSA, InternImage Multi-Task Pretraining over 300M 2024 Jun Dynamic OFA DOFA [83] Multiple Varied Global ViT MIM - 2024 Mar Multi-Modules UPetu [18] GeoSense - Global ConvNeXt-B SSL 0.65M 2024 May MP-MAE MMEarth [51] Multi-modal, Geospatial Data - Global ConvNeXt V2 MP-MAE 3.7M to 650M
    TABLE IV: Summary of foundation models for semantic segmentation in remote sensing, detailing architecture, pre-training datasets, resolution, geographic coverage, visual encoders, pretraining methods, and number of parameters.
    ! Year-Month Architecture Model Name Pre-training Dataset Resolution (m) Geographic Coverage visual encoder (backbone) Pretrain methods # of Params 2021 Oct ResNet-50 GeoKR [39] Levir-KR [39] 0.8 to 16 Global ResNet-50/VGG16 Geographical Knowledge Supervision 23.5M/138M 2021 Oct ResNet-50 SeCo [46] Sentinel-2 Imagery 10, 20, 60 200k Locations Worldwide ResNet-50 CL 23.5M 2022 Aug ResNet-50 GeCO [40] Levir-KR [39] 0.8 to 16 Global ResNet-50 SSL 23.5M 2022 Jul ViT/Swin Transformer RingMo [61] 2 Million RS Images 0.3 to 30 6 Continents ViT/Swin Transformer MIM - 2022 Nov ViT RVSA [71] MillionAID [44, 43] 0.5 to 153 Global ViT-B/ViTAE-B MAE 100M 2022 May ViTAEv2-S RSP [69] MillionAID [44, 43] 0.5 to 153 Global ViTAEv2-S/ ResNet-50/Swin-T Supervised Learning 24.8M/23.5M/29M 2022 Nov ViT-B CSPT [87] ImageNet-1K - Global ViT-B SSL 86M 2024 May ViT BFM [7] MillionAID [44, 43] 0.5 to 153 Global ViT MAE 86M/605.26M/1.36B/2.42B 2023 Apr TOV TOV [63] TOV-NI, TOV-RS - Global ViT SSL - 2023 Apr Teacher-student Self-distillation CMID [49] MillionAID [44, 43] Varied Global ResNet-50/Swin-B SSL 25.6M/87.8M 2023 Sep Multi-Branch RingMo-Sense [85] RS Spatiotemporal Dataset - Global Video Swim Transformer SSL - 2024 Feb MAE-based Framework Presto [66] Presto-21.5M 10 Global Transformer-based MAE 402K 2024 May ViT CtxMIM [86] WorldView-3 Imagery Varied Asia Swim-B MIM 88M 2023 Sep CNN-Transformer RingMo-lite [76] AID [81] 0.3 to 30 Global Swim Transformer + CNN FD-MIM 60% less than RingMo 2023 Oct MSFE+MMFH Feng et al. [21] Multi-modal Dataset Varied Global Transformer SSL - 2024 Jan ViT USat [31] Satlas Varied Global ViT MAE - 2024 March ViT FoMo-Bench [5] Multiple Varied Global ViT MAE 101M/110M 2024 Jan Teacher-Student Network GeRSP [29] ImageNet, MillionAID [44, 43] 0.5 to 153 Global ResNet-50 SSL, SL - 2024 Jan Generative ConvNet SMLFR [17] GeoSense 0.05 to 150 Multiple Continents ConvNeXt SSL 88M/197M 2024 May HiViT SARATR-X [37] SAR Datasets 0.1 to 3 Global HiViT MIM 66M 2024 Mar Factorized Multi-Modal Spatiotemporal Encoder SkySense [24] Multiple Varied Global Multiple CL 2.06B 2024 May Shared Encoder+ Task-Specific Decoders MTP [70] SAMRS Varied Global RVSA, InternImage Multi-Task Pretraining over 300M 2024 Jun Dynamic OFA DOFA [83] Multiple Varied Global ViT MIM -
    TABLE V: Summary of foundation models for object detection in remote sensing, detailing architecture, pre-training datasets, resolution, geographic coverage, visual encoders, pretraining methods, and number of parameters.
    ! Year-Month Architecture Model Name Pre-training Dataset Resolution (m) Geographic Coverage Visual encoder Pretrain methods # of Params 2021 Oct ResNet-50 SeCo[46] Sentinel-2 imagery 10, 20, 60 200k Locations Worldwide ResNet-50 CL 23.5M 2022 Jul ViT/Swin Transformer RingMo[61] 2M RS images 0.3 to 30 6 Continents ViT/Swin Transformer MIM - 2022 May ViT RSP[69] MillionAID 0.5 to 153 Global ViTAEv2-S/ ResNet-50/Swin-T Supervised Learning 24.8M/23.5M/29M 2021 Dec ResNet-34 MATTER[1] Sentinel-2 imagery - Rural and Remote Regions with Little Change ResNet34 SSL 21.3M 2023 Apr Teacher-Student Self-distillation CMID[49] MillionAID Varied Global ResNet-50/Swin-B SSL 25.6M/87.8M 2023 Jun CACo CACo[47] Sentinel-2 imagery 10 Global ResNet-18/50 SSL 11.7M/23.5M 2023 Aug Teacher-Student GFM[48] GeoPile - Global Swim Transformer Continual Pretraining - 2024 Apr DINO DINO-MC[77] SeCo-100K 10 to 60 Global ViT, Swim Transformer, ResNet, WRN SSL - 2023 Sep CNN-Transformer RingMo-lite[76] AID 0.3 to 30 Global Swim Transformer, CNN FD-MIM 60% less than RingMo 2023 Oct MSFE+MMFH Feng et al.[21] Multi-modal Dataset Varied Global Transformer SSL - 2024 Jan Dual-Branch SwiMDiff[65] Sen12MS Varied Global ResNet-18 SSL 11.7M 2024 Feb 3D GPT SpectralGPT[28] Sentinel-2 Imagery Varied Global ViT MAE 100M/300M/600M 2024 Mar Factorized Multi-Modal Spatiotemporal Encoder SkySense[24] Multiple Varied Global Multiple CL 2.06B 2024 May Shared Encoder+ Task-Specific Decoders MTP[70] SAMRS Varied Global RVSA, InternImage Multi-Task Learning over 300M
    TABLE VI: Summary of foundation models for change detection in remote sensing, detailing architecture, pre-training datasets, resolution, geographic coverage, visual encoders, pretraining methods, and number of parameters.
V-B Domain-Specific Tasks
Domain-specific tasks in remote sensing refer to applications that address particular real-world challenges by leveraging the capabilities of foundation models. These tasks include environmental monitoring, archaeology, agriculture and forestry, urban planning and development, and disaster management. To highlight the versatility of foundation models in remote sensing, we present figure 5 that categorizes models based on their applicability to various domain-specific tasks, as well as computer vision tasks. This figure serves as a quick reference for researchers to identify suitable models for their specific needs.
Figure 5: This diagram illustrates various domain-specific tasks in computer vision and their connections. Key areas include Environmental Monitoring, Agriculture and Forestry, Urban Planning and Development, Disaster Management, and Archaeology. Each domain comprises specific tasks such as change detection and object Detection. The relationships between these tasks and their application domains are depicted through connecting arrows, emphasizing the interconnected nature of computer vision applications across different fields.
  1. 1.
    Environmental Monitoring:
    Environmental monitoring involves using remote sensing models to track changes in the environment, such as deforestation, desertification, and pollution. These models help in understanding the impact of human activities and natural phenomena on the environment.
  2. 2.
    Archaeology:
    In archaeology, remote sensing models have been used to identify and analyze archaeological features and sites. These models help in detecting features such as ruins, artifacts, and ancient structures from satellite imagery. CNNs and ViTs are often utilized for their ability to process high-resolution images and capture fine details.
  3. 3.
    Agriculture and Forestry:
    In agriculture, remote sensing models are used to monitor crop health, estimate yields, and manage agricultural practices. These models help in optimizing resource use and improving agricultural productivity.
  4. 4.
    Urban Planning and Development:
    In urban planning and development, remote sensing models are used to monitor and analyze urban expansion, infrastructure development, and land use changes. These models help in managing urban growth, planning new developments, and assessing the impact of urbanization.
  5. 5.
    Disaster Management:
    Remote sensing models play a crucial role in disaster management by providing timely information on affected areas. These models are used to detect and assess damage from natural disasters such as earthquakes, floods, and hurricanes.
VI Discussion
The rapid advancement in foundation models for remote sensing underscore their transformative potential across various applications. As the field continues to evolve, it is crucial to synthesize the findings, address technical challenges, understand practical implications, and identify future research directions. In this section, we make a comprehensive analysis of these aspects, aiming to offer insights and guidance for future development and application of remote sensing foundation models.
VI-A Synthesis of Findings
In our survey of foundation models for remote sensing, we identified significant advancements and trends that highlights the evolving capabilities and applications of these models. The performance metrics of various models across different downstream tasks such as scene classification, semantic segmentation, object detection, and change detection reveal the following key findings:
VI-A1 Model Performance.
  • Classification. The performance of foundation models on the BigEarthNet dataset [60] for classification tasks shows variations in accuracy, as presented in table VIII. Overall, msGFM [25] has the top performance of 92.90% (mAP), followed closely by SkySense [24] with a performance of 92.09%. Other notable performers include DeCUR [72], which achieved an mAP of 89.70%, and DINO-MC [77], with an mAP of 88.75%. SeCo [46] also demonstrated strong performance with an mAP of 87.81%, while DINO-MM [73] reached an mAP of 87.10%. On the other hand, models like CACo [47] and FoMo-Bench [5] have mAP of 74.98% and F1-Score of 68.33% respectively, showing competitiveness but room for improvement.
    The high mAP scores of msGFM [25] and SkySense [24] highlight their efficiency in classifcation tasks, making them suitable for applications requiring high accuracy. Other foundation models, such as DINO-MM [73] and DeCUR [72], also provide strong performance with potential for further optimization. The variety in performance metrics underscores the evolving capabilities and specialization of foundation models in handling complex classification tasks within datasets like BigEarthNet [60].
  • Segmentation. For the segmentation tasks, we compared 12 foundation models which have been tested on the ISPRS Potsdam dataset. As shown in table VIII, SkySense [24] has the better performance out of all 12 models, with a mF1 Score of 93.99%. CMID [49] stands out with the highest mIoU of 87.04%, demonstrating its superior capability in accurately segmenting different regions within the dataset. For Overall Accuracy performance, BFM [7] has the highest OA score of 91.82%. Cross-Scale MAE [62], UPetu [18] and RSP [69] have mIoU scores of 76.17%, 83.17% and 65.30% respectively, showing competitive segmentation capabilities. GeoKR [39] reaches an mIoU of 70.48%, indicating robust segmentation performance but with room for improvement compared to CMID [49]. TOV scores the lowest mIoU at 60.34%, suggesting it may struggle with finer segmentation tasks compared to the other models.
    The performance metrics for the models applied to the ISPRS Potsdam dataset reveal significant variations in their effectiveness in segmentation tasks. SkySense [24] and CMID [49] emerge as top performers in mF1 score and mIoU, respectively, while SMLFR [17], RingMo [61], and RingMo-lite [76] demonstrate strong overall accuracy. These insights can guide the selection and optimization of models for specific remote sensing applications, ensuring the best possible performance for the task at hand.
    Dataset Model Performance (%) Metrics
    SeCo [46] 87.81 mAP
    CMC-RSSR [59] 82.90 mAP
    DINO-MM [73] 87.10 mAP
    CACo [47] 74.98 mAP
    GFM [48] 86.30 mAP
    DINO-MC [77] 88.75 mAP
    CROMA [22] 86.46 mAP
    DeCUR [72] 89.70 mAP
    CtxMIM [86] 86.88 mAP
    FG-MAE [75] 78.00 mAP
    USat [31] 85.82 mAP
    FoMo-Bench [5] 69.33 F1 Score
    SwiMDiff [65] 81.10 mAP
    SpectralGPT [28] 88.22 mAP
    SatMAE++ [52] 85.11 mAP
    msGFM [25] 92.90 mAP
    SkySense [24] 92.09 mAP
    BigEarthNet MMEarth [51] 78.6 mAP
    TABLE VII: This table provides an overview of the performance metrics for various models applied to the BigEarthNet dataset [60] for classification tasks. The performance is measured using mAP (Mean Average Precision) and F1 Score.
    Dataset Model Performance (%) Metrics
    GeoKR [39] 70.48 mIoU
    RSP [69] 65.30 mIoU
    RingMo [61] 91.74 OA
    RVSA [71] 91.22 OA
    TOV [63] 60.34 mIoU
    CMID [49] 87.04 mIoU
    RingMo-lite [76] 90.96 OA
    Cross-Scale MAE [62] 76.17 mIoU
    SMLFR [17] 91.82 OA
    SkySense [24] 93.99 mF1
    UPetu [18] 83.17 mIoU
    ISPRS Potsdam BFM [7] 92.58 OA
    TABLE VIII: This table provides an overview of the performance metrics for various models applied to the ISPRS Potsdam[30] dataset for segmentation tasks. The performance is measured using Mean Intersection over Union (mIoU) and Overall Accuracy (OA).
  • Object Detection. In table X, the performance of foundation models on the DOTA, DIOR, and DIOR-R datasets are evaluated based on their Mean Average Precision (mAP) and Average Precision at 50% (AP50).
    On the DOTA dataset, RVSA [71] achieves the highest mAP of 81.24% in accurately detecting objects, followed by SMLFR [17] and RSP [69] with mAP of 79.33% and 77.72%. CMID [49], GeRSP [29], and BFM [7] also demonstrate moderate performance with mAPs of 72.12%, 67.40% and 58.69%. For DIOR and DIOR-R dataset, MTP [70] and SkySense [24] are the top performers with an AP50 of 78% and an mAP of 78.73% respectively, showcasing their superior object detection capabilities. These insights can guide the selection and optimization of models to ensure the best possible performance for specific remote sensing applications.
  • Change Detection. For the change detection tasks, we compared the performance of foundation models on the OSCD and LEVIR-CD datasets. The models were evaluated based on their F1 Scores, which provide a balanced measure of precision and recall. As shown in the table, the performance varies significantly across different models and datasets.
    SkySense [24] achieves the highest F1 Score of 60.06% on the OSCD dataset, demonstrating its superior ability to accurately detect changes. GFM [48] follows with an F1 Score of 59.82%, indicating strong performance in change detection tasks. SpectralGPT [28] also performs well with an F1 Score of 54.29%. Other notable models include DINO-MC [77] with an F1 Score of 52.71% and CACo [47] with an F1 Score of 52.11%. SeCo [46] records the lowest F1 Score at 46.94%, suggesting it may require further optimization to enhance its change detection capabilities.
    In contrast, the LEVIR-CD dataset reveals higher performance metrics across the models. MTP [70] achieves the highest F1 Score of 92.67% and SkySense [24] follows closely with an F1 Score of 92.58%, demonstrating their robust performance. SWiMDiff reaches a lower F1 Score of 80.90% compared to its peers but still indicates effective performance in the LEVIR-CD dataset.
    Dataset Model Performance (%) Metrics
    RSP [69] 77.72 mAP
    RVSA [71] 81.24 mAP
    TOV [63] 26.10 mAP50
    CMID [49] 72.12 mAP
    GeRSP [29] 67.40 mAP
    SMLFR [17] 79.33 mAP
    DOTA BFM [7] 58.69 mAP
    RingMo [61] 75.80 mAP
    CSPT [87] 69.80 mAP
    RingMo-lite [76] 73.40 mAP
    GeRSP [29] 72.20 mAP
    DIOR MTP [70] 78.00 AP50
    RVSA [71] 71.05 mAP
    SMLFR [17] 72.33 mAP
    SkySense [24] 78.73 mAP
    MTP [70] 74.54 mAP
    DIOR-R BFM [7] 73.62 mAP
    TABLE IX: This table provides an overview of the performance metrics for various models applied to the DOTA[16, 79, 15] dataset, DIOR[38] and DIOR-R[9] dataset for object detection task. The performance is mainly measured using Mean Average Precision (mAP)
    Dataset Model Performance (%) Metrics
    SeCo [46] 46.94 F1 Score
    MATTER [1] 49.48 F1 Score
    CACo [47] 52.11 F1 Score
    GFM [48] 59.82 F1 Score
    SWiMDiff [65] 49.60 F1 Score
    SpectralGPT [28] 54.29 F1 Score
    SkySense [24] 60.06 F1 Score
    DINO-MC [77] 52.71 F1 Score
    OSCD MTP [70] 53.36 F1 Score
    RSP [69] 90.93 F1 Score
    RingMo [61] 91.86 F1 Score
    RIngMo-lite [76] 91.56 F1 Score
    SwiMDiff [65] 80.90 F1 Score
    SkySense [24] 92.58 F1 Score
    UPetu [18] 88.50 F1 Score
    LEVIR-CD MTP [70] 92.67 F1 Score
    TABLE X: This table provides an overview of the performance metrics for various models applied to the Onera Satellite Change Detection (OSCD) dataset[6] and the LEVIR-CD dataset[8] for change detection downstream tasks. The performance is measured using F1-Score.
  • Influence of Pre-training Methods. Various pre-training methods have a substantial impact on the performance of foundation models in remote sensing. Models pre-trained using self-supervised learning (SSL) techniques, such as contrastive learning (CL) and masked autoencoders (MAE), consistently exhibit superior performance compared to those pre-trained with traditional supervised learning. For instance, models like SkySense [24] and SeCo [46], which employ contrastive learning, achieve higher accuracy and robustness in tasks like scene classification and object detection. Additionally, models utilizing predictive coding techniques, such as SatMAE [11] and Scale-MAE [55], demonstrate enhanced capability in handling multi-temporal and multi-spectral data, leading to significant improvements in change detection and environmental monitoring applications. These findings highlight the critical role of innovative pre-training methods in maximizing the effectiveness of foundation models and suggest that continued exploration and refinement of these techniques are essential for advancing remote sensing capabilities.
VI-A2 Practical Implications
The advancements in foundation models have significant practical implications across various domains:
  • Environmental Monitoring. Models like GASSL [3] and SatMAE [11] offer detailed assessments of environmental changes, aiding in conservation efforts and policy-making. These models excel in monitoring deforestation, desertification, and pollution levels, providing actionable insights for environmental management. By integrating multi-spectral and temporal data, these models can track changes over time, allowing for early detection of environmental degradation and the formulation of timely interventions. This capability is particularly important for the sustainable management of natural resources, as well as reducing the impacts of climate change.
  • Agriculture and Forestry. Foundation models such as EarthPT [58] and GeCo [40] delivers valuable insights into crop health, yield predictions, and land use management, optimizing agricultural practices and resource allocation. For instance, RSP [69], leveraging multi-spectral data, enhances precision agriculture by accurately monitoring crop conditions and predicting yields. These models can detect early signs of crop stress, diseases, and pest infestations, enabling farmers to take proactive measures. Additionally, they aid in forestry management by providing detailed maps of forest cover, biomass estimation, and monitoring deforestation activities, thereby supporting conservation efforts and sustainable forestry practices.
  • Archaeology. The use of foundation models in archaeology revolutionizes the way archaeological features and sites are discovered, mapped, and analyzed. Models, such as GeoKR [39], RingMo [61], etc., can process high-resolution satellite imagery and multi-spectral data to enhance the detection and mapping of archaeological features that might be difficult to discern with the naked eye. Others like MATTER [1] can accomplish texture and material analysis to help identify various surface. They enable large-scale surveys, allowing archaeologists to identify potential sites of interest over vast areas efficiently. Although thorough exploration still requires on-site visits and excavations or other terrestrial investigations, these significantly improve the initial identification and mapping process. Additionally, these models can track changes over time, helping archaeologists monitor environmental and human impacts, providing crucial information for preservation and restoration. This enhances the efficiency and accuracy of surveys and opens new possibilities for discovering unknown sites.
  • Urban Planning and Development. Remote sensing models like CMID [49] and SkySense [24] are pivotal for monitoring urban expansion, infrastructure development, and land use changes. These models facilitate sustainable urban growth and development planning by providing high-resolution data analysis and trend forecasting. They enable city planners to assess the impact of urbanization on natural habitats, optimize land use, and plan infrastructure projects more effectively.
  • Disaster Management. Models such as OFA-Net[84], DOFA [83] and Prithvi [33] are instrumental in flood mapping, as well as fire detection. These models provide critical real-time data that helps in identifying affected areas quickly, enabling timely and effective response measures. This capability supports emergency responds in prioritizing resource allocation and implementing evacuation plans, thereby reducing the impact of natural disasters. Additionally, these models assist in post-disaster recovery by assessing damage and monitoring the recovery process over time. By integrating various data sources, they enhance the ability to make informed decisions, coordinate response efforts, and plan for future disaster mitigation strategies.
VI-B Future Direction
Future research should prioritize several key areas:
  • Efficient Model Development: Exploring techniques such as model distillation, pruning, and quantization to reduce computational requirements without compromising performance is crucial. Additionally, developing scalable architectures that efficiently handle ultra-high-resolution images is essential. For instance, applying pruning techniques to models like SatMAE [11] could maintain performance while reducing computational load.
  • Multi-Modal Data Integration: Enhancing methods for integrating and processing multi-modal data (e.g., combining optical and radar imagery) will provide more comprehensive insights. Research on advanced SSL techniques capable of leveraging multi-modal data is necessary. The OFA-Net [84] framework, which integrates multi-model data, serves as a promising direction for future models to emulate and improve upon.
  • Interdisciplinary Collaboration: Promoting collaboration between remote sensing experts, AI researchers, and domain specialists can address complex challenges and drive innovation. For example, partnerships between AI researchers and environmental scientists can refine models like GASSL [3] for better environmental monitoring and conservation efforts.
VI-C Limitation
This survey has several limitations:
  • Scope and Coverage. The review focuses on foundation models released between June 2021 and June 2024. While the scope of this review is extensive and covers many significant developments, it is not exhaustive. Some recent advancements and innovations in the field may not be included due to their release timing or the lack of sufficient evaluation metrics at the time of writing. Consequently, certain cutting-edge models that have emerged in the latter part of this period or that have not yet been thoroughly evaluated might be omitted. This limitation underscores the need for readers to seek out the most current research and updates beyond the scope of this survey.
  • Evolving Field. The field of AI and remote sensing is rapidly evolving, with continuous advancements and breakthroughs occurring at a fast pace. This dynamic nature necessitates ongoing reviews and updates to ensure the relevance and comprehensiveness of the survey. New techniques, methodologies, and models are constantly being developed, which can significantly impact the state of the art. Therefore, it is essential to recognize that this survey represents a snapshot in time and that continuous monitoring of the literature is required to capture the latest advancements and emerging trends. This approach will help maintain an up-to-date understanding of the field and incorporate new findings as they become available.
VII Conclusion
In this comprehensive survey, we have reviewed the advancements in foundation models for remote sensing developed between June 2021 and June 2024. We categorized these models into vision and vision-language types, highlighting their unique methodologies and capabilities.
Our analysis covered various advanced techniques, including self-supervised learning (SSL), Vision Transformers (ViTs), and residual neural networks (ResNets). These models have significantly improved performance in tasks such as scene classification, semantic segmentation, and object detection, as well as in domain-specific applications like environmental monitoring, digital archaeology, agriculture, urban planning, and disaster management.
While significant progress has been made, several challenges persist, such as the need for more diverse and high-quality datasets, high computational requirements, and task-specific difficulties. Addressing these challenges will require further research and collaboration across disciplines.
In summary, this survey provides a detailed overview of the current state of foundation models in remote sensing, offering valuable insights and identifying future research directions. We recommend continued efforts in developing efficient model architectures, enhancing multi-modal data integration, and expanding dataset diversity to fully realize the potential of these models in remote sensing.
Acknowledgements. This work was supported by Vanderbilt Seeding Success Grant, Vanderbilt Discovery Grant, and VISR Seed Grant. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant. This works was also supported by NSF NAIRR Pilot Award NAIRR240055. This manuscript has been co-authored by ORNL, operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S.Department of Energy. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.